# Two Characterizations of Geometrically Infinite Actions on Gromov Hyperbolic Spaces

### v1

@@ 2.2. Geometrically finite actions
For a hyperbolic element $g\in G$, the translation length $\tau(g)$ of $g$ is defined by
$\tau(g):=\inf_{x\in X}d(x,gx)$
and is strictly positive. The translation length is invariant under conjugacy. We may also define the stable translation length by $\tau^{\infty}(g):=\lim\limits_{n\to\infty}\frac{1}{n}d(o,g^{n}o)$.
Lemma 2.8.
[CDP06, Proposition 10. 6.4] Let $g$ be a hyperbolic isometry of a $\delta$-hyperbolic space $X$. Then
$\lvert\tau(g)-\tau^{\infty}(g)\rvert\leq 16\delta.$
We denote its limit points on the boundary by $g^{+}:=\lim\limits_{n\to\infty}g^{n}o\in\partial X$ and $g^{-}:=\lim\limits_{n\to\infty}g^{-n}o\in\partial X$. These are the two fixed points of $g$, and an axis of $g$ is a geodesic line $g^{+},g^{-}]$. By Lemma 2.7, there exists a constant $D_{0}>0$ depending only on $\delta$ such that the Hausdorff distance between any two axes of $g$ is bounded by $D_{0}$.
Lemma 2.9.
There exists a constant $D_{1}>0$ such that
$\tau(g)\leq d(x,gx)\leq\tau(g)+D_{1}$
for any hyperbolic element $g\in G$ and any point $x$ on some axis of $g$.
Proof.
Suppose $\gamma$ is an axis of $g$ and $x\in\gamma$. By definition we have $d(x,gx)\geq\tau(g)$. Note that $g^{n}x\in g^{n}\gamma$ and $d_{\text{Haus}}(g^{n}\gamma,\gamma)\leq D_{0}$ for any $n\in\mathbb{Z}$. Let $x_{n}\in\pi_{\gamma}(g^{n}x)$ for every $n\in\mathbb{Z}$. Then for every $n\in\mathbb{N}$ we have
$d(x_{n+1},x_{n})\geq d(g^{n+1}x,g^{n}x)-d(g^{n+1}x,x_{n+1})-d(g^{n}x,x_{n})=d(x,gx)-2D_{0}.$
Hence
$d(x_{n},x)=\sum\limits_{i=1}^{n}d(x_{i-1},x_{i})\geq n(d(x,gx)-2D_{0}).$
This implies that $\tau^{\infty}(g)\geq d(x,gx)-2D_{0}$. Therefore, by Lemma 2.8 we have
$\tau(g)\geq\tau^{\infty}(g)-16\delta\geq d(x,gx)-2D_{0}-16\delta.$
Setting $D_{1}=2D_{0}+16\delta$ completes the proof. ∎
@@
We define the weak-convex hull $\mathrm{CH}(\Lambda G)$ of $\Lambda G$ as the union of all bi-infinite geodesics with two endpoints in $\Lambda G$. By $\delta$-thin triangle property, one sees that $\mathrm{CH}(\Lambda G)$ is a $G$-invariant uniformly quasi-convex subset of $X$. Namely, there exists a constant $\sigma$ (depending only on $\delta$) so that any geodesic with endpoints in $\mathrm{CH}(\Lambda G)$ is contained in the $\sigma$-neighborhood of $\mathrm{CH}(\Lambda G)$. Further, the topological boundary of $\mathrm{CH}(\Lambda G)$ in $\partial X$ coincides with the limit set $\Lambda G$.
Following [BOW12], we say that a family of quasi-convex subsets in $X$ is sufficiently separated if it is $r$-separated for a sufficiently large constant $r$ depending on $\delta$. The action of $G$ on $X$ is called geometrically finite if there exists a sufficiently separated $G$-invariant family $\{B_{p}:p\in\Pi\}$ of horoballs in $X$, where each horoball $B_{p}$ is centered at a point $p$ in a $G$-invariant set $\Pi\subset\partial X$, and the action of $G$ on the complement $\mathrm{CH}(\Lambda G)\setminus\bigcup_{p\in\Pi}B_{p}$ is cocompact.
It is a fundamental fact that for proper actions on hyperbolic spaces, the geometric definition of geometric finiteness given above is equivalent to the dynamical definition in terms of the boundary action. We record this equivalence in the following lemma.
@@ 3. Escaping Closed Geodesics
In this section, we assume that a group $G$ acts properly by isometries on a proper $\delta$-hyperbolic space $X$ for some $\delta\geq 0$. Moreover, we assume that the action is non-elementary. We fix a basepoint $o\in X$. Recall that $d_{\text{Haus}}(\gamma_{1},\gamma_{2})\leq D_{0}$ for any hyperbolic element $g\in G$ and any axes $\gamma_{1},\gamma_{2}$ of $g$.
The goal of this section is to establish Theorem 1.1. The nontrivial direction is that if there is no escaping sequence of hyperbolic elements, then the action is geometrically finite. Accordingly, we assume throughout that there exists $R_{0}>0$ such that $d(\gamma,Go)\leq R_{0}$ for every hyperbolic element $g\in G$ and some axis $\gamma$ of $g$.
Together with the standing assumption that the action is non-elementary (which ensures the existence of independent hyperbolic elements), we have a finiteness statement about conjugacy classes.
@@
This yields $d(o,h^{-1}gh\,o)=d(ho,gho)<2R_{0}+D_{1}+r$. Since the action of $G$ on $X$ is proper, the set $\{k\in G:d(o,ko)<2R_{0}+D_{1}+r\}$ is finite. Therefore, the set of conjugates $h^{-1}gh$ belongs to this finite set. Consequently, there are only finitely many distinct conjugacy classes of such elements $g$. ∎
@@ 3.1. The main proposition
Assume that the non-conical limit set $\Lambda^{nc}G$ is non-empty; otherwise there is nothing to prove. Our next goal is to prove the following key estimate, which describes the geometry of lines connecting a non-conical limit point to its translates.
@@
There exists a constant $R_{1}=R_{1}(R_{0},\delta)$ such that $d([\xi,g\xi],Go)\leq R_{1}$ for any $\xi\in\Lambda^{\mathrm{nc}}G$ and any hyperbolic element $g\in G$ with $\tau(g)\geq 2D_{0}+200\delta$.
@@
For any hyperbolic element $g\in G$ with $\tau(g)\geq 2D_{0}+200\delta$, our standing hypothesis ensures the existence of an element $h\in G$, an axis $\gamma$ of $g$ and a point $x\in\gamma$ such that $d(x,ho)\leq R_{0}$. Then for any $n\in\mathbb{Z}$, the points $g^{n}ho\in Go$ and $g^{n}x\in g^{n}\gamma$ also satisfy $d(g^{n}ho,g^{n}x)\leq R_{0}$ and $g^{n}\gamma$ is also an axis of $g$. Hence $d_{\text{Haus}}(g^{n}\gamma,\gamma)\leq D_{0}$ and $d(g^{n}ho,\gamma)\leq R_{0}+D_{0}$ for every $n\in\mathbb{Z}$.
@@
Choose points $y\in\pi_{\gamma}(\xi)$ and $z\in\pi_{\gamma}(g\xi)$. Then we have $gy\in\pi_{g\gamma}(g\xi)$. Since $d_{\text{Haus}}(\gamma,g\gamma)\leq D_{0}$, Lemma 2.6 implies $d(z,gy)\leq 2D_{0}+100\delta$. Hence
@@
Let $[g]$ denote the conjugacy class of a hyperbolic element $g$ in $G$. Then there exists a constant $R_{[g]}\geq 0$ such that $d([\xi,g\xi],Go)\leq R_{[g]}$ for any $\xi\in\Lambda^{\mathrm{nc}}G$ and any element $g\in[g]$.
@@
There exists a constant $R_{2}\geq R_{0}$ such that $d([\xi,g\xi],Go)\leq R_{2}$ for any $\xi\in\Lambda^{\mathrm{nc}}G$ and any hyperbolic element $g\in G$.
@@
Figure 1. Illustration for the proof of Proposition 3.2.
@@
Choose an integer $N$ such that $N>R_{2}+2\delta$ and $g_{N}\xi\neq\eta$. We claim that the product $g_{n}g_{N}$ is a hyperbolic element for all sufficiently large $n$.
Choose neighborhoods $U$ and $V$ of $\xi$ in $\partial X$ such that $\xi\in V\subset\overline{V}\subsetneq U$ and $\eta\notin g_{N}\overline{U}$. Such neighborhoods exist because $\partial X$ is a compact Hausdorff space with no isolated points and $g_{N}\xi\neq\eta$. Since $g_{n}x\to\xi$ locally uniformly on $\overline{X}\setminus\{\eta\}$, for sufficiently large $n$ we have $g_{n}(g_{N}\overline{U})\subseteq V\subsetneq U$. By Lemma 2.1, this implies that $g_{n}g_{N}$ is hyperbolic, proving the claim.
Now let $n$ be large enough so that $g_{n}g_{N}$ is hyperbolic and $n>R_{2}+2\delta$. By Lemma 3.5, we have $d([\xi,g_{n}g_{N}\xi],Go)\leq R_{2}$. Since the ideal triangle $\triangle(\xi,g_{n}\xi,g_{n}g_{N}\xi)$ is $2\delta$-thin, the geodesic $\xi,g_{n}g_{N}\xi]$ lies within $2\delta$ of $\xi,g_{n}\xi]\cup[g_{n}\xi,g_{n}g_{N}\xi]$. Therefore,
@@ 3.2. Identifying the cusp regions
We now introduce the key geometric sets that will be used to analyze the dynamics near a non-conical limit point. These sets are essential for establishing the geometric finiteness criterion in the sufficiency direction of Theorem 1.1. Our strategy is to construct, for each non-conical limit point $\xi$, a geometric object $B_{\xi,R}$ that plays the role of an invariant horoball. We will show that for sufficiently large $R$, the translates of these sets are pairwise disjoint, and that their truncations $S_{\xi,R}=B_{\xi,R}\cap Y_{R}$ have compact quotient under the stabilizer $G_{\xi}$. This will prove that every $\xi\in\Lambda^{nc}G$ is a bounded parabolic point, thereby completing the argument.
Recall that the constant $R_{0}$ satisfies $d(\gamma,Go)\leq R_{0}$ for every hyperbolic element $g\in G$ and some axis $\gamma$ of $g$. For $R\geq R_{0}$ and $\xi\in\Lambda^{nc}G$, define
$Y_{R}$ $=N_{R}(Go)=\{x\in X:d(x,Go)\leq R\},$
[…]
The following lemma confirms the equivariance of this construction.
@@
Observe that $x\in B_{\xi,R}$ if and only if there exists a geodesic ray $\gamma=[x,\xi]$ lying outside $Y_{R}$. This is equivalent to $g\gamma=[gx,g\xi]$ lying outside $gY_{R}=Y_{R}$, which means $gx\in B_{g\xi,R}$. ∎
@@
For any $\xi\in\Lambda^{nc}G$ and any $g\in G$ with $g\xi\neq\xi$, we have $B_{\xi,R}\cap B_{g\xi,R}=\varnothing$ for all $R>R_{\xi}+2\delta$.
@@
Suppose $\xi\in\Lambda^{nc}G$ and $R\geq R_{0}$. Then any geodesic ray $\gamma$ with $\gamma(\infty)=\xi$ is eventually contained in $B_{\xi,R}$; that is, there exists $a\geq 0$ such that $\gamma([a,\infty))\subset B_{\xi,R}$.
@@
Suppose $\xi\in\Lambda^{nc}G$ and $R\geq R_{\xi}+2\delta$. Then $B_{\xi,R}$ is a closed subset in $X$.
@@
Let $\{x_{n}\}$ be a sequence in $B_{\xi,R}$ converging to a point $x\in X$. Then there exists a sequence $\{\gamma_{n}=[x_{n},\xi]\}$ of geodesic rays such that $d(\gamma_{n},Go)\geq R$. By the Arzelà–Ascoli lemma, there exists a subsequence $\{n_{i}\}$ such that $\gamma_{n_{i}}$ converges locally uniformly to a geodesic ray $\gamma=[x,\xi]$. Then $d(\gamma,Go)\geq R$, which implies $x\in B_{\xi,R}$. ∎
@@
For any $\xi\in\Lambda^{nc}G$, $g\in G$ with $\xi\neq g\xi$, and any $R>R_{\xi}+2\delta$, we have $S_{\xi,R}\cap S_{g\xi,R}=\varnothing$.
@@
Assume the action of $G$ on $X$ is geometrically finite. By Lemma 2.10, there exists a $G$-invariant family of pairwise disjoint horoballs $\{B_{p}:p\in\Lambda^{nc}G\}$ such that the action of $G$ on the complement $\mathrm{CH}(\Lambda G)\setminus\bigcup_{p\in\Lambda^{nc}G}B_{p}$ is cocompact.
By the pairwise disjointness of the horoballs, any axis of any hyperbolic element $g\in G$ must intersect the complement $X\setminus\bigcup_{p\in\Lambda^{nc}G}B_{p}$. Suppose the subspace $\mathrm{CH}(\Lambda G)$ is $D$-quasiconvex for some $D>0$. Then every axis $\gamma$ of $g$ lies in the $D$-neighborhood of $CH(\Lambda G)$. Hence $\gamma\cap N_{D}(\mathrm{CH}(\Lambda G))\setminus\bigcup_{p\in\Lambda^{nc}G}B_{p}\neq\varnothing$.
@@
Conversely, assume there exists a constant $R_{0}>0$ such that $d(\gamma,Go)\leq R_{0}$ for every hyperbolic element $g$ in $G$ and some axis $\gamma$ of $g$. We must show that the action is geometrically finite. By Lemma 2.10, it is enough to prove that every non-conical limit point is a bounded parabolic point.
Fix a non-conical limit point $\xi\in\Lambda^{nc}G$ and choose $R>R_{\xi}+2\delta$, where $R_{\xi}$ is the constant from Proposition 3.2. Consider the quotient map $\pi:Y_{R}\to Y_{R}/G$ induced by the proper action. The image $\pi(S_{\xi,R})$ is a closed subset of the compact space $Y_{R}/G$, hence $\pi(S_{\xi,R})$ is compact. For any $g\in G$ with $g\xi\neq\xi$, Lemma 3.6 gives $gS_{\xi,R}=S_{g\xi,R}$, and Corollary 3.10 ensures that $S_{\xi,R}\cap S_{g\xi,R}=\varnothing$. Consequently, the $\pi$-preimage of $\pi(S_{\xi,R})$ is precisely the $G_{\xi}$-orbit of $S_{\xi,R}$, inducing a homeomorphism $\pi(S_{\xi,R})\cong S_{\xi,R}/G_{\xi}$, where $G_{\xi}$ denotes the stabilizer of $\xi$ in $G$. The compactness of $\pi(S_{\xi,R})$ thus implies that the action of $G_{\xi}$ on $S_{\xi,R}$ is cocompact.
@@
To prove $\xi$ is bounded parabolic, we construct a compact fundamental domain for the action of $G_{\xi}$ on $\partial X\setminus\{\xi\}$. Let $K\subset S_{\xi,R}$ be a compact fundamental domain for the cocompact action of $G_{\xi}$ on $S_{\xi,R}$ (such a $K$ exists by the argument above). Define a subset $D\subset\partial X\setminus\{\xi\}$ by
@@
The existence of such a compact fundamental domain $\overline{D}$ means precisely that $\xi$ is a bounded parabolic point. Since $\xi\in\Lambda^{nc}G$ was arbitrary, every non-conical limit point is bounded parabolic. Applying Lemma 2.10 (or the standard dynamical characterization of geometric finiteness), we conclude that the action of $G$ on $\Lambda G$ is geometrically finite.

### v2

@@ 2.3. Geometrically finite actions
Definition 2.21.
The weak-convex hull $\mathrm{CH}(\Lambda G)$ of $\Lambda G$ is the union of all bi-infinite geodesics with two endpoints in $\Lambda G$.
By the $\delta$-thin triangle property, one sees that $\mathrm{CH}(\Lambda G)$ is a $G$-invariant uniformly quasi-convex subset of $X$. Namely, there exists a constant $D$ (depending only on $\delta$) such that any geodesic with endpoints in $\mathrm{CH}(\Lambda G)$ is contained in the $D$-neighborhood of $\mathrm{CH}(\Lambda G)$. Further, the topological boundary of $\mathrm{CH}(\Lambda G)$ in $\partial X$ coincides with the limit set $\Lambda G$.
Following [BOW12], we say that a family of quasi-convex subsets in $X$ is sufficiently separated if it is $r$-separated for a sufficiently large constant $r$ depending on $\delta$. The following notion was based on the work of Marden [MAR74] for Kleinian groups, and further developed by Bowditch in various contexts [BOW93, BOW95, BOW12].
Definition 2.22.
The action of $G$ on $X$ is called geometrically finite if there exists a sufficiently separated $G$-invariant family $\{B_{p}:p\in\Pi\}$ of horoballs in $X$, where each horoball $B_{p}$ is centered at a point $p$ in a $G$-invariant set $\Pi\subset\partial X$, and the action of $G$ on the complement $\mathrm{CH}(\Lambda G)\setminus\bigcup_{p\in\Pi}B_{p}$ is cocompact.
It is a fundamental fact that for proper actions on hyperbolic spaces, the geometric definition of geometric finiteness is equivalent to the dynamical definition in terms of the boundary action in Definition 2.5. We record this equivalence in the following lemma.
@@ 3. Escaping Closed Geodesics
In this section, we assume that a group $G$ acts properly by isometries on a proper $\delta$-hyperbolic space $X$ for some $\delta\geq 0$. Moreover, we assume that the action is non-elementary. We fix a basepoint $o\in X$. Recall that $d_{\mathrm{Haus}}(\gamma_{1},\gamma_{2})\leq D_{0}$ for any hyperbolic element $g\in G$ and any axes $\gamma_{1},\gamma_{2}$ of $g$.
The goal of this section is to establish Theorem 1.1.
The nontrivial direction is to show that, if there is no escaping sequence of hyperbolic elements, then the action is geometrically finite. Accordingly, we impose the following assumption up to Subsection 3.3, where this implication, together with the converse direction in Theorem 1.1, will be proved using the preparatory results established below.
Non-Escaping Hyperbolic Elements (NEHE) Assumption
There exists $R_{0}>0$ such that for every hyperbolic element $g\in G$, some axis $\gamma$ of $g$ satisfies $d(\gamma,Go)\leq R_{0}$.
Since the action is non-elementary, there are infinitely many conjugacy classes of hyperbolic elements. An immediate consequence of the NEHE Assumption is the following finiteness statement about conjugacy classes.
@@
This yields $d(o,h^{-1}gh\,o)=d(ho,gho)<2R_{0}+D_{1}+r$. Since the action of $G$ on $X$ is proper, the set $\{k\in G:d(o,ko)<2R_{0}+D_{1}+r\}$ is finite. Therefore, there is an element $h^{-1}gh$ which is conjugate to $g$ and belongs to this finite set. Consequently, there are only finitely many conjugacy classes of such elements. ∎
@@ 3.1. The main proposition
Assume that the non-conical limit set $\Lambda^{nc}G$ is non-empty; otherwise there is nothing to prove. Maintaining NEHE Assumption, our next goal is to prove the following key estimate, which describes the geometry of lines connecting a non-conical limit point to its translates.
@@
There exists a constant $R_{1}=R_{1}(R_{0},\delta)$ such that $d([\xi,g\xi],Go)\leq R_{1}$ for any $\xi\in\Lambda^{\mathrm{nc}}G$ and for any hyperbolic element $g\in G$ with $\tau(g)\geq 2D_{0}+200\delta$.
@@
For any hyperbolic element $g\in G$ with $\tau(g)\geq 2D_{0}+200\delta$, our NEHE Assumption ensures the existence of an element $h\in G$, an axis $\gamma$ of $g$ and a point $x\in\gamma$ such that $d(x,ho)\leq R_{0}$. Then for any $n\in\mathbb{Z}$, the points $g^{n}ho\in Go$ and $g^{n}x\in g^{n}\gamma$ also satisfy $d(g^{n}ho,g^{n}x)\leq R_{0}$ and $g^{n}\gamma$ is also an axis of $g$. Hence $d_{\mathrm{Haus}}(g^{n}\gamma,\gamma)\leq D_{0}$ and $d(g^{n}ho,\gamma)\leq R_{0}+D_{0}$ for every $n\in\mathbb{Z}$.
@@
Choose points $y\in\pi_{\gamma}(\xi)$ and $z\in\pi_{\gamma}(g\xi)$. Then we have $gy\in\pi_{g\gamma}(g\xi)$. Since $d_{\mathrm{Haus}}(\gamma,g\gamma)\leq D_{0}$, Lemma 2.20 implies $d(z,gy)\leq 2D_{0}+100\delta$. Hence
@@
Let $[g]$ denote the conjugacy class of a hyperbolic element $g$ in $G$. Then there exists a constant $R_{[g]}\geq 0$ such that $d([\xi,g\xi],Go)\leq R_{[g]}$ for any $\xi\in\Lambda^{\mathrm{nc}}G$ and for any element $g\in[g]$.
@@
There exists a constant $R_{2}\geq R_{0}$ such that $d([\xi,g\xi],Go)\leq R_{2}$ for any $\xi\in\Lambda^{\mathrm{nc}}G$ and for any hyperbolic element $g\in G$.
@@
Figure 1. Illustrating the hyperbolic elements $g_{n}g_{N}$ in the proof of Proposition 3.2.
@@
Fix an integer $N$ such that $N>R_{2}+2\delta$ and $g_{N}\xi\neq\eta$. We claim that the product $g_{n}g_{N}$ is a hyperbolic element for all sufficiently large $n$.
Choose neighborhoods $U$ and $V$ of $\xi$ in $\partial X$ such that $\xi\in V\subset\overline{V}\subsetneq U$ and $\eta\notin g_{N}\overline{U}$. Such neighborhoods exist because $\partial X$ is a compact Hausdorff space with no isolated points and $g_{N}\xi\neq\eta$. Since $g_{n}x\to\xi$ locally uniformly on $\overline{X}\setminus\{\eta\}$, for sufficiently large $n$ we have $g_{n}(g_{N}\overline{U})\subseteq V\subsetneq U$. See Figure 1. By Lemma 2.3, this implies that $g_{n}g_{N}$ is hyperbolic, proving the claim.
Now let $n$ be large enough so that $g_{n}g_{N}$ is hyperbolic and $n>R_{2}+2\delta$. By Lemma 3.5, we have $d([\xi,g_{n}g_{N}\xi],Go)\leq R_{2}$. Since the ideal triangle $\triangle(\xi,g_{n}\xi,g_{n}g_{N}\xi)$ is $2\delta$-thin, the geodesic $\xi,g_{n}g_{N}\xi]$ lies within $2\delta$ of $\xi,g_{n}\xi]\cup[g_{n}\xi,g_{n}g_{N}\xi]$. See Figure 2. Therefore,
@@ 3.2. Identifying the cusp regions
We now introduce the key geometric sets that will be used to analyze the dynamics near a non-conical limit point. These sets are essential for establishing the geometric finiteness criterion in the sufficiency direction of Theorem 1.1.
Our strategy is to construct, for each non-conical limit point $\xi$, a geometric object $B_{\xi,R}$, which plays the role of an invariant horoball. We will show that for sufficiently large $R$, the translates of these sets are pairwise disjoint, and that their truncations $S_{\xi,R}=B_{\xi,R}\cap Y_{R}$ have compact quotient under the stabilizer $G_{\xi}$. This will prove that every $\xi\in\Lambda^{nc}G$ is a bounded parabolic point, thereby completing the argument.
Recall the NEHE Assumption that the constant $R_{0}$ satisfies $d(\gamma,Go)\leq R_{0}$ for every hyperbolic element $g\in G$ and some axis $\gamma$ of $g$.
For $R\geq R_{0}$ and $\xi\in\Lambda^{nc}G$, define $Y_{R}:=\{x\in X:d(x,Go)\leq R\}$ and
$B_{\xi,R}:=\left\{x\in X\;\middle|\;\begin{aligned} &\text{there exists a geodesic ray }\gamma=[x,\xi]\\ &\text{such that }d(\gamma,Go)\geq R\end{aligned}\right\}.$
The following lemma justifies the equivariance of this construction.
@@
Observe that $x\in B_{\xi,R}$ if and only if there exists a geodesic ray $\gamma=[x,\xi]$ lying outside $Y_{R}$. This is equivalent to that $g\gamma=[gx,g\xi]$ lies outside $gY_{R}=Y_{R}$, which means $gx\in B_{g\xi,R}$. ∎
@@
For any $\xi\in\Lambda^{nc}G$ and for any $g\in G$ with $g\xi\neq\xi$, we have $B_{\xi,R}\cap B_{g\xi,R}=\varnothing$ for all $R>R_{\xi}+2\delta$.
@@
Let $\xi\in\Lambda^{nc}G$ and $R\geq R_{0}$. Then any geodesic ray $\gamma$ with $\gamma(\infty)=\xi$ is eventually contained in $B_{\xi,R}$; that is, there exists $a\geq 0$ such that $\gamma([a,\infty))\subset B_{\xi,R}$.
@@
Let $\xi\in\Lambda^{nc}G$ and $R\geq R_{\xi}+2\delta$. Then $B_{\xi,R}$ is a closed subset in $X$.
@@
Let $\{x_{n}\}$ be a sequence of points in $B_{\xi,R}$ converging to a point $x\in X$. Then there exists a sequence $\{\gamma_{n}=[x_{n},\xi]\}$ of geodesic rays such that $d(\gamma_{n},Go)\geq R$. By the Arzelà–Ascoli lemma, there exists a subsequence $\{n_{i}\}$ such that $\gamma_{n_{i}}$ converges locally uniformly to a geodesic ray $\gamma=[x,\xi]$. Then $d(\gamma,Go)\geq R$, which implies $x\in B_{\xi,R}$. Thus, $B_{\xi,R}$ is closed. ∎
@@
For any $\xi\in\Lambda^{nc}G$, $g\in G$ with $\xi\neq g\xi$, and for any $R>R_{\xi}+2\delta$, we have $S_{\xi,R}\cap S_{g\xi,R}=\varnothing$.
@@
This direction is well-known and we include a proof for completeness. Assume the action of $G$ on $X$ is geometrically finite. By Lemma 2.23, there exists a $G$-invariant family of pairwise disjoint horoballs $\{B_{p}:p\in\Lambda^{nc}G\}$ such that the action of $G$ on the complement $\mathrm{CH}(\Lambda G)\setminus\bigcup_{p\in\Lambda^{nc}G}B_{p}$ is cocompact.
It is a known fact that no bi-infinite geodesics are contained in a horoball. Since horoballs $\{B_{p}:p\in\Lambda^{nc}G\}$ are pairwise disjoint, any axis of any hyperbolic element $g\in G$ must intersect the complement $X\setminus\bigcup_{p\in\Lambda^{nc}G}B_{p}$. Suppose the weak hull $\mathrm{CH}(\Lambda G)$ is $D$-quasiconvex for some $D>0$. Then every axis $\gamma$ of $g$ lies in the $D$-neighborhood of $\mathrm{CH}(\Lambda G)$. Hence $\gamma\cap N_{D}(\mathrm{CH}(\Lambda G))\setminus\bigcup_{p\in\Lambda^{nc}G}B_{p}\neq\varnothing$.
@@
Conversely, assume there exists a constant $R_{0}>0$ such that $d(\gamma,Go)\leq R_{0}$ for every hyperbolic element $g$ in $G$ and some axis $\gamma$ of $g$ (i.e. NEHE assumption). We must show that the action is geometrically finite. By Lemma 2.23, it is enough to prove that every non-conical limit point is a bounded parabolic point.
Fix a non-conical limit point $\xi\in\Lambda^{nc}G$ and choose $R>R_{\xi}+2\delta$, where $R_{\xi}$ is the constant from Proposition 3.2. Consider the quotient map
$\pi:Y_{R}\to Y_{R}/G$
induced by the proper action. The image $\pi(S_{\xi,R})$ is a closed subset of the compact space $Y_{R}/G$, hence $\pi(S_{\xi,R})$ is compact.
For any $g\in G$ with $g\xi\neq\xi$, Lemma 3.6 gives $gS_{\xi,R}=S_{g\xi,R}$, and Corollary 3.10 ensures that $S_{\xi,R}\cap S_{g\xi,R}=\varnothing$. Consequently, the $\pi$-preimage of $\pi(S_{\xi,R})$ is precisely the $G_{\xi}$-orbit of $S_{\xi,R}$, inducing a homeomorphism
$\pi(S_{\xi,R})\cong S_{\xi,R}/G_{\xi}$
where $G_{\xi}$ denotes the stabilizer of $\xi$ in $G$. The compactness of $\pi(S_{\xi,R})$ thus implies that the action of $G_{\xi}$ on $S_{\xi,R}$ is cocompact.
@@
It remains to prove $\xi$ is bounded parabolic. To that end, we construct a compact fundamental domain for the action of $G_{\xi}$ on $\partial X\setminus\{\xi\}$. First, let $K\subset S_{\xi,R}$ be a compact fundamental domain for the cocompact action of $G_{\xi}$ on $S_{\xi,R}$. Define a boundary subset $D\subset\partial X\setminus\{\xi\}$ by
@@
The existence of such a compact fundamental domain $\overline{D}$ means precisely that $\xi$ is a bounded parabolic point. Since $\xi\in\Lambda^{nc}G$ was arbitrary, every non-conical limit point is bounded parabolic. Applying Lemma 2.23 (or the standard dynamical characterization of geometric finiteness), we conclude that the action of $G$ on $\Lambda G$ is geometrically finite. The Sufficiency direction direction follows and Theorem 1.1 is proved.